# Dense-SfM: Structure from Motion with Dense Consistent Matching

### v1

@@ Abstract
We present Dense-SfM, a novel Structure from Motion (SfM) framework designed for dense and accurate 3D reconstruction from multi-view images. Sparse keypoint matching, which traditional SfM methods often rely on, limits both accuracy and point density, especially in texture-less areas. Dense-SfM addresses this limitation by integrating dense matching with a Gaussian Splatting (GS) based track extension which gives more consistent, longer feature tracks. To further improve reconstruction accuracy, Dense-SfM is equipped with a multi-view kernelized matching module leveraging transformer and Gaussian Process architectures, for robust track refinement across multi-views. Evaluations on the ETH3D and Texture-Poor SfM datasets show that Dense-SfM offers significant improvements in accuracy and density over state-of-the-art methods.
@@ 1 Introduction
Figure 1: Comparison of SfM results between quantized-matching and our matching. To solve the fragmentary track problem,  [24] uses quantized matching, which can compromise the accuracy and completeness of 3D points due to quantization. In contrast, our Dense-SfM pipeline leverages Gaussian Splatting to solve the problem, offering more accurate and dense 3D points. Thanks to dense matching, our pipeline can also be applied to texture-less objects, further improving the quality of 3D points and pose estimation.
@@
Structure from Motion (SfM) is a fundamental task in computer vision that enables the reconstruction of 3D structures from multi-view images of a scene. It has been a widely researched problem, focusing on reconstructing accurate camera poses and 3D point cloud of a scene, benefiting a variety of downstream tasks such as visual localization [48, 45, 50, 59, 18], AR/VR [73, 42], multi-view stereo [19, 54, 52] and neural rendering [36, 27, 75]. Over the years, several well-established methods have been developed to address the SfM pipeline [70, 51, 56, 39].
@@
Recently, there has been an increasing demand for dense and accurate 3D reconstructions in various 3D vision tasks, including dense reconstruction [67, 16], object pose estimation [58, 23, 4] and neural rendering [27, 75]. However, existing detector-based methods have limitations in achieving sufficient accuracy and density of 3D points, especially in texture-less regions where obtaining matchable keypoints is challenging, limiting the performance of downstream tasks. With the advent of deep learning and attention mechanism [64], semi-dense and dense matching methods [57, 68, 5, 13, 14] have been introduced, enabling the reliable matches even in low-texture areas and producing denser match results than classical detector-based methods. This advancement has led to improved performance across many tasks, including homography and relative pose estimation [2, 7, 29]. However, due to their pairwise matching approach, the matching methods usually yield fragmented tracks [^1] A keypoint is assigned a set of 2D coordinates, called feature track or track in short, on neighbor images where the keypoint is matched. as mentioned in [24], making direct integration with existing SfM pipelines challenging. To solve this inconsistency issue,  [24] applied quantization to the matched keypoints, and then refined the multi-view tracks to obtain accurate camera pose and point cloud. While the quantization strategy can increase track length and consistency, enhancing the multi-view refinement results, it can significantly limit the number of matches and accuracy, reducing the quality of the resulting SfM model, as exemplified in Fig. 1.
@@
In this paper, we propose a novel SfM framework that leverages dense matching while addressing the limitations of pairwise matching methods. As shown in Fig. 1, our approach reconstructs a dense and accurate 3D point cloud without relying on quantized matching, avoiding the inherent limitations. Specifically, our framework first performs dense matching using a dense matcher, e.g., DKM [13] or RoMa [14], to construct an initial SfM model, where most tracks are initially fragmentary, i.e., of short (mostly a pair of images) lengths. Next, we initialize small 3D Gaussians based on the initial SfM’s point cloud, integrating attributes like color and position. These Gaussians are then optimized and densified through Gaussian Splatting (GS) [27, 75], filling in the scene while keeping the initial Gaussians’ (SfM points’) positions fixed. We then infer the visibility of each initial SfM point per image by rendering through the ray from camera poses. In case that the 3D point is visible on the image, we include the image into the track of the 3D point thereby increasing the track length, which is called track extension. Note that track extension enables us to resolve the problem of short tracks in pairwise matching. Lastly, we perform our proposed SfM refinement pipeline, which includes a multi-view kernelized matching module and geometric bundle adjustment, to further optimize the camera poses and 3D points. Our track refinement module is mainly built on two architectures, transformer and Gaussian Process [43], which incorporate both feature-based and positional information, resulting in notable performance gains over existing refinement modules as demonstrated in our ablation study (Tab. 3).
@@
We evaluate our Dense-SfM framework on the ETH3D and Texture-Poor SfM datasets, where it outperforms state-of-the-art SfM systems across multiple metrics. Thanks to the dense matching coupled with track extension via Gaussian Splatting, as well as our novel track refinement module, our framework achieves accurate camera pose estimations and dense 3D reconstructions.
@@
  • A track extension method that, based on Gaussian Splatting, helps find new visible images thereby increasing track length of 3D point.
@@ 2 Related Work
Establishing 2D correspondences through feature matching is a fundamental component for Structure-from-Motion (SfM) and visual SLAM. Traditional feature matching pipelines typically involve: i) detecting sparse keypoints [47, 31, 46, 8, 44, 10], ii) computing visual descriptors [33, 63, 61, 40, 12, 34], and iii) performing nearest-neighbor search or using learned matchers [49, 74, 17]. While this approach is efficient and computationally fast, especially with sparse points, it struggles in challenging scenarios such as low-textured regions, where keypoint detection is unreliable, or repetitive patterns, where matching can fail—leading to suboptimal SfM results.
@@
However, semi-dense and dense matching often results in fragmentary feature tracks, making it difficult to apply on SfM pipeline [55]. To solve this, [24] introduced quantization matching that merges nearby subpixel matches into grid nodes, improving track consistency. Yet, this approach can compromise the accuracy of the initial camera poses and point clouds, heavily relying on refinement modules and reducing point cloud density.
In contrast, our framework does not apply quantization but leverages Gaussian Splatting to assess visibility of 3D points. By projecting 3D points onto new images on which the points are visible, we obtain longer and consistent tracks without sacrificing the initial matching accuracy, which leads to improved refinement performance and, finally, better SfM results.
@@
Feature correspondence-based SfM methods have been extensively studied [37, 3, 15]. Incremental SfM [1, 56, 70] has become a widely adopted approach, with COLMAP [51] recognized as one of the most popular frameworks. This method relies on 2D feature matching verified via two-view epipolar geometry or homography [21] between image pairs. A small set of images is chosen to initialize the reconstruction, followed by the sequential registration of additional images. For each new image, the PnP problem [32] is solved using the matching results, after which new 3D points are triangulated and refined through bundle adjustment [62].
In addition to feature matching, several deep learning approaches have been introduced to improve bundle adjustment. [38] proposes the process of 2D point tracks with permutation-equivariant linear encoder layers to directly regress 3D point coordinates and camera poses. [60] introduces graph attention networks to estimate camera parameters and 3D coordinates of keypoints tracked across multiple views. Similarly, [66] leverages a probabilistic diffusion framework to replicate the iterative nature of bundle adjustment, minimizing epipolar error by utilizing 2D matches and enforcing geometric constraints across image pairs.
@@
To address this, several previous methods have proposed multi-view keypoint refinement models based on flow [11] or dense features [30], which lead to significant improvements in SfM accuracy.  [11] refines keypoint locations using a geometric cost constrained by local flow between two views, followed by global optimization using a multi-view graph of relative feature geometry constraints. Similarly,  [30] refines keypoint locations through feature-metric keypoint adjustment by optimizing a direct cost over discriminative dense feature maps along the tracks.
The current state-of-the-art SfM refinement method, [24], introduces a multi-view matching module that iteratively refines the tracks obtained from the COLMAP framework. The module extracts local patches from each image and feeds them into a transformer-based network to refine the keypoint’s location by enforcing multi-view consistency.
@@ 3.2 Track Extension via Gaussian Splatting
In the initial SfM, most of the tracks are short, i.e. consisting of mostly two images, due to the fragmentary nature of feature tracks in pair-wise matching, which can limit the performance of track refinement, as shown in [24].
@@
To determine the visibility of each 3D point from a camera pose, we apply a visibility filter using Gaussian Splatting (GS) [27, 75], which supports fast training and rendering. We initialize a GS model with the 3D points obtained in the initial SfM. Specifically, we treat them as small 3D Gaussians setting each mean $\boldsymbol{\mu}\in\mathbb{R}^{3}$ to the 3D point coordinates, rotation $R$ to the identity matrix, opacity $\alpha$ to 1, and scale $S$ based on the policy from [26], which uses focal length and depth information. These Gaussians are optimized and densified to cover the entire scene, while keeping the initial Gaussian’s parameters fixed. The details of initialization and training process is described in supplementary materials.
After training the GS model, we utilize it when evaluating the visibility $M$ of all SfM points (represented as small Gaussians) from each camera pose using the following formula:
@@
where $R$ is the set of rays from the camera center (of a camera pose) for rendering the image, $\alpha_{\text{SfM}}$ is the opacity of Gaussian representing the SfM point (set to $1$), $N_{SfM}$ the number of Gaussians along the ray $r$ from the camera center to the SfM point, and $\alpha_{j}$ the opacity of $j$-th intermediate Gaussian. If a visibility score is larger than $\epsilon_{\textit{v}}$, we consider the SfM point, i.e. $N_{SfM}$-th Gaussian, visible from the view of the camera pose.
For each 3D point, we extend its track by adding its projected 2D point(s) onto the visible view(s). To do that, we project a 3D point $P_{j}$ onto the visible image using camera pose $\{R_{i},t_{i}\}$ and intrinsic matrix $C_{i}$ as follows:
@@ 3.3.1 Feature Track Refinement
Following [24], we split a track into the keypoint of a reference view and those of query views (views from the track except the reference view) with a policy of median depth value. Then, we feed image patches centered at each keypoint into a CNN feature extractor, obtaining feature maps ${\mathbf{F}_{R},\mathbf{F}_{Q_{i}}\in\mathbb{R}^{p\times p\times c}}$, where $\mathbf{F}_{R}$ is from the reference view and $\mathbf{F}_{Q_{i}}$’s are from query views. As shown in Fig. 4, our multi-view kernelized matching architecture takes as input these feature maps of reference and query views.
Our proposed matching architecture consists of two paths, feature-based and coordinate-embedding-based ones. On the feature-based path (denoted with Multi-view Transformer in Fig. 4), we utilize a local feature transformer to obtain the transformed feature map $\hat{\mathbf{F}}_{R}$ and $\hat{\mathbf{F}}_{Q_{i}}$’s through self and cross-attention between $\mathbf{F}_{R}$ and $\mathbf{F}_{Q}$’s. On the coordinate-embedding-based path (denoted with Gaussian-Process in Fig 4), we employ a Gaussian Process [43] to utilize the positional information, deriving a coordinate embedding feature $\mu(\mathbf{F}_{R}|\mathbf{F}_{Q_{i}})$ for each query view. Following  [13], we use an exponential cosine similarity kernel:
@@
Fig. 5 illustrates how the refined track is obtained. First, we obtain the refined coordinate of keypoint on the reference view. From the output of multi-view kernelized matching, $S_{Q_{i}}$’s, we estimate the confidence value of each pixel as the candidate of refined coordinate on the reference view. Then, as shown in Fig. 5 (top), we select, as the refined coordinate on the reference view, the pixel with the highest value, i.e. the highest confidence from $\sum_{i}S_{Q_{i}}$.
@@
We set up training as in [24] using MegaDepth dataset [29]. We use image bags sampled from the SfM model provided in the dataset, in which ground truth tracks are derived from the given camera pose and MVS depth map. For training, we perturb the ground truth 2D locations of tracks in the query views with random noise, providing these as input to our refinement module. Unlike [24], which relies on statistics (variance) as uncertainty measure, we directly learn confidence score through a neural network. Consequently, our loss function is not simply formulated as an $L_{2}$ loss to minimize the distance between the refined tracks and the ground truth tracks. Instead, we jointly predict and learn the confidence $S_{Q_{i}}>0$ for each query view, and include it in the regression loss as follows:
@@
where $n_{t}$ is the number of feature tracks, $n_{j}$ the number of query views in a track $j$, $N$ the total number of refined keypoints on all the query views, $p_{Q_{i}}$ and $p_{gt}$ the learned 2D coordinate from $Q_{i}$ (illustrated in Fig. 5) and the ground truth 2D coordinate, respectively, and $s_{Q_{i}}$ the confidence score of prediction $p_{Q_{i}}$, which is obtained, from $S_{Q_{i}}$, at the pixel of the reference view’s ground truth keypoint. The second term in the loss is used to avoid being under-confident, where we set $\alpha$ as 20.

### v2

@@ Abstract
We present Dense-SfM, a novel Structure from Motion (SfM) framework designed for dense and accurate 3D reconstruction from multi-view images. Sparse keypoint matching, which traditional SfM methods often rely on, limits both accuracy and point density, especially in texture-less areas. Dense-SfM addresses this limitation by integrating dense matching with a Gaussian Splatting (GS) based track extension which gives more consistent, longer feature tracks. To further improve reconstruction accuracy, Dense-SfM is equipped with a multi-view kernelized matching module leveraging transformer and Gaussian Process architectures, for robust track refinement across multi-views. Evaluations on the ETH3D and Texture-Poor SfM datasets show that Dense-SfM offers significant improvements in accuracy and density over state-of-the-art methods. Project page: https://icetea-cv.github.io/densesfm/.
@@ 1 Introduction
Figure 1: Comparison of SfM results between quantized-matching and our matching. To solve the fragmentary track problem, DFSfM [25] uses quantized matching, which can compromise the accuracy and completeness of 3D points due to quantization. In contrast, our Dense-SfM pipeline leverages Gaussian Splatting to solve the problem, offering more accurate and dense 3D points. Thanks to dense matching, our pipeline can also be applied to texture-less objects, further improving the quality of 3D points and pose estimation.
@@
Structure from Motion (SfM) is a fundamental task in computer vision that enables the reconstruction of 3D structures from multi-view images of a scene. It has been a widely researched problem, focusing on reconstructing accurate camera poses and 3D point cloud of a scene, benefiting a variety of downstream tasks such as visual localization [49, 46, 51, 61, 20], AR/VR [75, 42], multi-view stereo [21, 55, 53] and neural rendering [37, 28, 77]. Over the years, several well-established methods have been developed to address the SfM pipeline [72, 52, 58, 39, 43].
@@
Recently, there has been an increasing demand for dense and accurate 3D reconstructions in various 3D vision tasks, including dense reconstruction [68, 18], object pose estimation [60, 24, 5] and neural rendering [28, 77]. However, existing detector-based methods have limitations in achieving sufficient accuracy and density of 3D points, especially in texture-less regions where obtaining matchable keypoints is challenging, limiting the performance of downstream tasks. With the advent of deep learning and attention mechanism [65], semi-dense and dense matching methods [59, 69, 6, 15, 16] have been introduced, enabling the reliable matches even in low-texture areas and producing denser match results than classical detector-based methods. This advancement has led to improved performance across many tasks, including homography and relative pose estimation [2, 8, 31]. However, due to their pairwise matching approach, the matching methods usually yield fragmented tracks [^1] A keypoint is assigned a set of 2D coordinates, called feature track or track in short, on neighbor images where the keypoint is matched. as mentioned in Detector-Free SfM (DFSfM) [25], making direct integration with existing SfM pipelines challenging. To solve this inconsistency issue, DFSfM applied quantization to the matched keypoints, and then refined the multi-view tracks to obtain accurate camera pose and point cloud. While the quantization strategy can increase track length and consistency, enhancing the multi-view refinement results, it can significantly limit the number of matches and accuracy, reducing the quality of the resulting SfM model, as exemplified in Fig. 1.
@@
In this paper, we propose a novel SfM framework that leverages dense matching while addressing the limitations of pairwise matching methods. As shown in Fig. 1, our approach reconstructs a dense and accurate 3D point cloud, avoiding the inherent limitations of quantized matching. Specifically, our framework first performs dense matching using a dense matcher, e.g., DKM [15] or RoMa [16], to construct an initial SfM model, where most tracks are initially fragmentary, i.e., of short (mostly a pair of images) lengths. Next, we initialize small 3D Gaussians based on the initial SfM’s point cloud, integrating attributes like color and position. These Gaussians are then optimized and densified through Gaussian Splatting (GS) [28, 77], filling in the scene while keeping the initial Gaussians’ (SfM points’) positions fixed. We then infer the visibility of each initial SfM point per image by rendering through the ray from camera poses. In case that the 3D point is visible on the image, we include the image into the track of the 3D point thereby increasing the track length—a process we refer to as Track Extension. Note that Track Extension enables us to resolve the problem of short tracks in pairwise matching. Lastly, we perform our proposed SfM refinement pipeline, which includes a multi-view kernelized matching module and geometric bundle adjustment, to further optimize the camera poses and 3D points. Our track refinement module is mainly built on two architectures, transformer and Gaussian Process [44], which incorporate both feature-based and positional information, resulting in notable performance gains over existing refinement modules as demonstrated in our ablation study (Tab. 3).
@@
We evaluate our Dense-SfM framework on the ETH3D, Texture-Poor SfM and Image Matching Challenge (IMC) 2021 datasets, where it outperforms state-of-the-art SfM systems across multiple metrics. Thanks to the dense matching coupled with track extension via Gaussian Splatting, as well as our novel track refinement module, our framework achieves accurate camera pose estimations and dense 3D reconstructions.
@@
  • A track extension method that, based on Gaussian Splatting, helps find new images on which 3D point is visible, thereby increasing track length of the 3D point.
@@ 2 Related Work
Establishing 2D correspondences through feature matching is a fundamental component for SfM and visual SLAM. Traditional feature matching typically involves i) detecting keypoints [48, 33, 47, 9, 45, 12], ii) computing descriptors [34, 64, 62, 40, 14, 35] and iii) performing descriptor matching [50, 76, 19]. Despite its efficiency, it struggles in challenging scenarios such as low-textured regions, where keypoint detection is unreliable, and repetitive patterns, where matching often fails—leading to suboptimal SfM results.
@@
However, semi-dense and dense matching often results in fragmentary feature tracks, making it difficult to apply on SfM pipeline [56]. To solve this, DFSfM [25] introduced quantization matching that merges nearby subpixel matches into grid nodes, improving track consistency. Yet, this approach can compromise the accuracy of the initial camera poses and point clouds, heavily relying on refinement modules and reducing point cloud density.
In contrast, our framework leverages Gaussian Splatting to assess visibility of 3D points. By projecting 3D points onto new images on which the points are visible, we obtain longer and consistent tracks without sacrificing the initial matching accuracy, which leads to improved refinement performance and, finally, better SfM results.
@@
Feature correspondence-based SfM methods have been extensively studied [38, 3, 17]. Incremental SfM [1, 58, 72] has become a widely adopted approach, with COLMAP [52] recognized as one of the most popular frameworks.
Recently, differentiable SfM methods [66, 67, 4, 57, 70, 11] have aimed to simplify or entirely replace the traditional SfM pipeline. VGGSfM [67] manages all SfM components as fully differentiable, enabling end-to-end training. DUSt3R [70] predicts pairwise pointmaps and refines them through optimization-based global alignment. Its successors, MASt3R [30] and MASt3R-SfM [11], improve pixel matching for localization by predicting additional per-pixel features and leveraging its encoder for fast image retrieval, leading to a newly proposed SfM pipeline.
@@
To address this, several previous methods have proposed multi-view keypoint refinement models based on flow [13] or dense features [32], which lead to significant improvements in SfM accuracy. PatchFlow [13] refines keypoint locations using a geometric cost constrained by local flow between two views, followed by global optimization using a multi-view graph of relative feature geometry constraints. Similarly, PixSfM [32] refines keypoint locations through feature-metric keypoint adjustment by optimizing a direct cost over discriminative dense feature maps along the tracks. DFSfM [25], introduces a multi-view matching module that iteratively refines the tracks obtained from the COLMAP framework. The module extracts local patches from each image and feeds them into a transformer-based network to refine the keypoint’s location by enforcing multi-view consistency.
@@ 3.2 Track Extension via Gaussian Splatting
In the initial SfM, most of the tracks are short, i.e. consisting of mostly two images, due to the fragmentary nature of feature tracks in pair-wise matching, which can limit the performance of track refinement, as shown in DFSfM [25].
@@
To determine the visibility of each 3D point from a camera pose, we apply a visibility filter using Gaussian Splatting (GS) [28, 77], which supports fast training and rendering. We initialize a GS model with the 3D points obtained in the initial SfM. Specifically, we treat them as small 3D Gaussians setting each mean $\boldsymbol{\mu}\in\mathbb{R}^{3}$ to the 3D point coordinates, rotation $R$ to the identity matrix, opacity $\alpha$ to 1, and scale $S$ based on the policy from Splatam [27], which uses focal length and depth information. These Gaussians are optimized and densified to cover the entire scene, while keeping the initial Gaussian’s parameters fixed. The details of initialization and training process is described in supplementary material.
After training the GS model, we utilize it for evaluating the visibility $M$ of all SfM points (represented as small Gaussians) from each camera pose using the following formula:
@@
where $R$ is the set of rays from the camera center (of a camera pose) for rendering the image, $\alpha_{\text{SfM}}$ is the opacity of Gaussian representing the SfM point (set to $1$), $N_{SfM}$ the number of Gaussians along the ray $r$ from the camera center to the SfM point, and $\alpha_{j}$ the opacity of $j$-th intermediate Gaussian. If the highest visibility score among the pixels rendered from the Gaussian representing the SfM point is larger than $\epsilon_{\textit{v}}$, we consider the SfM point, i.e. $N_{SfM}$-th Gaussian, visible from the view of the camera pose. For each 3D point, we extend its track by adding its projected 2D point(s) onto the visible view(s). To do that, we project a 3D point $P_{j}$ onto the visible image using camera pose $\{R_{i},t_{i}\}$ and intrinsic matrix $C_{i}$ as follows:
@@ 3.3.1 Feature Track Refinement
Following DFSfM, we split a track into the keypoint of a reference view and those of query views (views from the track except the reference view) with a policy of median depth value. Then, we feed image patches centered at each keypoint into a CNN feature extractor, obtaining feature maps ${\mathbf{F}_{R},\mathbf{F}_{Q_{i}}\in\mathbb{R}^{p\times p\times c}}$, where $\mathbf{F}_{R}$ is from the reference view and $\mathbf{F}_{Q_{i}}$’s are from query views. As shown in Fig. 4, our multi-view kernelized matching architecture takes as input these feature maps of reference and query views.
Our proposed matching architecture consists of two paths, feature-based and coordinate-embedding-based ones. On the feature-based path (denoted with Multi-view Transformer in Fig. 4), we utilize a local feature transformer to obtain the transformed feature map $\hat{\mathbf{F}}_{R}$ and $\hat{\mathbf{F}}_{Q_{i}}$’s through self and cross-attention between $\mathbf{F}_{R}$ and $\mathbf{F}_{Q}$’s. On the coordinate-embedding-based path (denoted with Gaussian-Process in Fig. 4), we employ a Gaussian Process [44] to utilize the positional information, deriving a coordinate embedding feature $\mu(\mathbf{F}_{R}|\mathbf{F}_{Q_{i}})$ for each query view. Following DKM [15], we use an exponential cosine similarity kernel:
@@
Fig. 5 illustrates how the refined track is obtained. First, we obtain the refined coordinate of keypoint on the reference view. From the output of multi-view kernelized matching, $\{S_{Q_{i}}\}$, we estimate the confidence value of each pixel as the candidate of refined coordinate on the reference view. As shown in Fig. 5 (top), we select, as the refined coordinate on the reference view, the pixel with the highest value, i.e. the highest confidence from $\sum_{i}S_{Q_{i}}$.
@@
We set up training as in DFSfM [25] using MegaDepth dataset [31]. We use image bags sampled from the SfM model provided in the dataset, in which ground truth tracks are derived from the given camera pose and MVS depth map. For training, we perturb the ground truth 2D locations of tracks in the query views with random noise, providing these as input to our refinement module. Unlike DFSfM, which relies on statistics (variance) as uncertainty measure, we directly learn confidence score through a neural network. Consequently, our loss function is not simply formulated as an $L_{2}$ loss to minimize the distance between the refined tracks and the ground truth tracks. Instead, we jointly predict and learn the confidence $S_{Q_{i}}>0$ for each query view, and include it in the regression loss as follows:
@@
where $n_{t}$ is the number of feature tracks, $n_{j}$ the number of query views in a track $j$, $N$ the total number of refined keypoints on all the query views, $p_{Q_{i}}$ and $p_{gt}$ the learned 2D coordinate from $Q_{i}$ (as illustrated in Fig. 5) and the ground truth 2D coordinate, respectively, and $s_{Q_{i}}$ the confidence score of prediction $p_{Q_{i}}$, which is obtained, from $S_{Q_{i}}$, at the pixel of the reference view’s ground truth keypoint. The second term in the loss is used to avoid being under-confident, where we set $\alpha$ as 20.